# Information About Other Agents in Mechanism Design

### v3

@@ 6 Conclusion
The main finding of the paper is that increasing the information agents have about the types of other agents can serve to eliminate undesired equilibria in full implementation problems. Consequently, there are social choice rules that are not originally fully implementable under a given information structure but become fully implementable after agents receive noisy signals about the types of other agents. We provide a sufficient condition that guarantees the existence of a more informative information structure that can allow an SCF $F$ that is not fully implementable under some full-support common prior to become fully implementable. We conclude by discussing some limitations of the analysis.
The results are dependent on the specifics of agents’ beliefs in different information structures. In particular, given a non-fully-implementable SCF, any increase in information agents have about the types of other agents would need to be common knowledge for there to potentially exist a mechanism that fully implements the SCF under the new information structure. While these assumptions are standard in the implementation literature, a recent strand of literature in mechanism design focuses on the robustness of mechanisms (pioneered by Bergemann and Morris (2005)), arguing that mechanism design should not rely excessively on common knowledge assumptions among the agents and the planner. The dependence on common knowledge assumptions is certainly a limitation on the results presented, but it is worth noting the main finding — that increasing information an agent has about other agents can eliminate undesired equilibria — does not depend on a single, specific change in information. Rather, in every example and theorem presented, the signal an agent receives that eliminates an undesired equilibrium is simply required to be above a certain level of accuracy. In all of the examples and results presented, whenever a signal $\sigma$ of accuracy $10.5<\tau<$ 1 eliminates an undesired equilibrium, any signal with the same realizations but higher accuracy $\tau^{\prime}>\tau$ also eliminates the equilibrium. Thus, while the agents do need to fully understand when other agents have received additional information, the exact accuracy of the signals received does not need to be known.
Appendix A Appendix A: A more robust example
In the example provided in section 4, agent 3 is indifferent between the truth-telling equilibrium and the deception $\alpha$ regardless of the social choice rule, raising the question of whether it is simply describing a knife-edge case. But the example can be modified to accommodate this concern.
There are still three agents $i=1,2,3$, where each agent $i$ has types $\theta_{i},\theta_{i}^{\prime},$ and 6 possible outcomes $A=\{a,b,c,d,e,f\}.$ The SCF $F$ remains the same, as follows.
$F\mid\theta_{3}$ $\theta_{2}$ $\theta_{2}^{\prime}$
$\theta_{1}$ $a$ $b$
$\theta_{1}^{\prime}$ $e$ $d$
$F\mid\theta_{3}^{\prime}$ $\theta_{2}$ $\theta_{2}^{\prime}$
$\theta_{1}$ $a$ $c$
$\theta_{1}^{\prime}$ $e$ $d$
Modify the agents’ preferences to get the following.
Type Preferences
$\theta_{1}$ $e\,(6)\succ a\,(4)\succ b,c\,(3)\succ d\,(2)\succ f\,(1)$
$\theta_{1}^{\prime}$ $b,c\,(6)\succ d\,(4)\succ e\,(3)\succ a\,(2)\succ f\,(1)$
$\theta_{2}$ $f\,(25)\succ b,c\,(7)\succ a\,(3)\succ e\,(2)\succ d\,(1)$
$\theta_{2}^{\prime}$ $f\,(20)\succ e\,(4.5)\succ d\,(3)\succ b,c\,(2)\succ a\,(1)$
$\theta_{3}$ $a\,(5)\succ b\,(4)\succ f\,(3)\succ c\,(2)\succ d,e\,(1)$
$\theta_{3}^{\prime}$ $a\,(5)\succ c\,(4)\succ f\,(3)\succ b\,(2)\succ d,e\,(1)$
Modify the common prior $\mu$ to be defined by
$\mu(\theta_{1},\theta_{2},\cdot)=1/10,\mu(\theta_{1},\theta_{2}^{\prime},\cdot)=3/10,\mu(\theta_{1}^{\prime},\theta_{2},\cdot)=9/20,\mu(\theta_{1}^{\prime},\theta_{2}^{\prime},\cdot)=3/20.$
Moreover, the type of agent 3 is independent of the types of agents 1 and 2, and $\mu(\cdot,\cdot,\theta_{3})=\mu(\cdot,\cdot,\theta_{3}^{\prime})=1/2$.
We follow similar steps as in the main example to show that $F$ is incentive-compatible and that the direct mechanism contains the equilibria $\alpha=((\theta_{1}^{\prime},\theta_{1}),(\theta_{2}^{\prime},\theta_{2}),(\theta_{3},\theta_{3}^{\prime}))$, $\alpha_{1}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3},\theta_{3}^{\prime})),\alpha_{2}=((\theta_{1},\theta_{1}),(\theta_{2}^{\prime},\theta_{2}^{\prime}),(\theta_{3},\theta_{3}^{\prime})),\alpha_{3}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3}^{\prime},\theta_{3}^{\prime})),\alpha_{4}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3}^{\prime},\theta_{3}))$, and $\alpha_{5}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3},\theta_{3}))$. It is also analogous to show that all of the $\alpha_{i}$ for $1\leq i\leq 5$ can be undermined by agents 1 and 2.
The deception $\alpha$ still violates Bayesian monotonicity under the specified common prior, meaning that no agent can undermine $\alpha$. As before, first consider agent 1. In order to define a $y:\Theta_{-1}\to A$ such that
$\mathbb{E}(u_{1}(y(\alpha_{-1}(\theta_{-1})),\theta_{1})\mid\theta_{1})>\mathbb{E}(u_{1}(F(\alpha_{-1}(\theta_{-1},\theta_{1})),\theta_{1})\mid\theta_{1})$
it must be that (1) $y(\theta_{2},\cdot)=e$, $y(\theta_{2}^{\prime},\cdot)\succeq_{\theta_{1}}d$, and at least one of $y(\theta_{2}^{\prime},\theta_{3})\succeq_{\theta_{1}}b,c$ or $y(\theta_{2}^{\prime},\theta_{3}^{\prime})\succeq_{\theta_{1}}b,c$ holds; (2) $y(\theta_{2},\theta_{3})=e$, $y(\theta_{2},\theta_{3}^{\prime})=a$ and $y(\theta_{2}^{\prime},\cdot)\succeq_{\theta_{1}}b,c$; or (3) $y(\theta_{2},\theta_{3})=a$, $y(\theta_{2},\theta_{3}^{\prime})=e$ and $y(\theta_{2}^{\prime},\cdot)\succeq_{\theta_{1}}b,c$. If (1) applies, compute that
$\mathbb{E}(u_{1}(y(\theta_{-1}),\theta_{1})\mid\theta_{1})\geq\frac{27}{8}>\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})=\frac{13}{4},$
which is not a valid alternate SCF $y.$ If any of the other cases apply, it must also be that
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})<\mathbb{E}(u_{1}(y(\theta_{-1}),\theta_{1})\mid\theta_{1}).$
By similar reasoning, in order to define a $y:\Theta_{-1}\to A$ such that
$\mathbb{E}(u_{1}(y(\alpha_{-1}(\theta_{-1})),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})>\mathbb{E}(u_{1}(F(\alpha(\theta_{-1},\theta_{1}^{\prime})),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})$
it must also be that
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})<\mathbb{E}(u_{1}(y(\theta_{-1}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime}).$
Thus agent 1 cannot undermine $\alpha$. For agent 2, the utility experienced from outcome $f$ is again too high for her to include it as a possible outcome in any alternate SCF $y,$ so agent 2 cannot undermine $\alpha$. Since agent 3 is no longer indifferent between $\alpha$ and truth-telling, it remains to show that no $y:\Theta_{-3}\to A$ exists that makes $F$ Bayesian monotonic. Note that
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}),\theta_{3})\mid\theta_{3})=\frac{1}{10}u_{3}(a,\theta_{3})+\frac{3}{10}u_{3}(b,\theta_{3})+\frac{9}{20}u_{3}(e,\theta_{3})+\frac{3}{20}u_{3}(d,\theta_{3})$
while
$\mathbb{E}(u_{3}(F(\alpha(\theta_{-3},\theta_{3})),\theta_{3})\mid\theta_{3})=\frac{3}{20}u_{3}(a,\theta_{3})+\frac{9}{20}u_{3}(b,\theta_{3})+\frac{3}{10}u_{3}(e,\theta_{3})+\frac{1}{10}u_{3}(d,\theta_{3}).$
Then there are 2 ways to define a $y:\Theta_{-3}\to A$ to get
$\mathbb{E}(u_{3}(y(\alpha(\theta_{-3})),\theta_{3})\mid\theta_{3})>\mathbb{E}(u_{3}(F(\alpha(\theta_{-3},\theta_{3})),\theta_{3})\mid\theta_{3}).$
An alternate SCF $y$ must either satisfy (1) $y(\theta_{1},\cdot)=a$ and $y(\theta_{1}^{\prime},\cdot)\in\{d,e\}$ or (2) $y(\theta_{1}^{\prime},\theta_{2})\succ_{\theta_{3}}d,e$ or $y(\theta_{1}^{\prime},\theta_{2}^{\prime})\succ_{\theta_{3}}d,e$. If (1) applies, it is clear that
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}),\theta_{3})\mid\theta_{3})<\mathbb{E}(u_{3}(y(\theta_{-3}),\theta_{3})\mid\theta_{3}).$
Now suppose (2) applies. Then in order for
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}),\theta_{3})\mid\theta_{3})\geq\mathbb{E}(u_{3}(y(\theta_{-3}),\theta_{3})\mid\theta_{3})$
it must be that
$u_{3}(a,\theta_{3})+3u_{3}(b,\theta_{3})-u_{3}(y(\theta_{1},\theta_{2}),\theta_{3})-3u_{3}(y(\theta_{1},\theta_{2}^{\prime}),\theta_{3})$
$\geq\frac{3}{2}\left(3u_{3}(y(\theta_{1}^{\prime},\theta_{2}),\theta_{3})+u_{3}(y(\theta_{1}^{\prime},\theta_{2}^{\prime}),\theta_{3})-3u_{3}(e,\theta_{3})-u_{3}(d,\theta_{3})\right).$
On the other hand, for
$\mathbb{E}(u_{3}(y(\alpha(\theta_{-3})),\theta_{3})\mid\theta_{3})>\mathbb{E}(u_{3}(F(\alpha(\theta_{-3},\theta_{3})),\theta_{3})\mid\theta_{3})$
to hold, it must be that
$3u_{3}(y(\theta_{1}^{\prime},\theta_{2}),\theta_{3})+u_{3}(y(\theta_{1}^{\prime},\theta_{2}^{\prime}),\theta_{3})-3u_{3}(e,\theta_{3})-u_{3}(d,\theta_{3})$
$>\frac{3}{2}(u_{3}(a,\theta_{3})+3u_{3}(b,\theta_{3})-u_{3}(y(\theta_{1},\theta_{2}),\theta_{3})-3u_{3}(y(\theta_{1},\theta_{2}^{\prime}),\theta_{3})).$
Since all quantities are positive, it is impossible for (iii) and (iv) to hold simultaneously. Thus there cannot be a $y:\Theta_{-3}\to A$ that allows $F$ to satisfy Bayesian monotonicity.
Similar reasoning as what was given in the main example implies the signal $\sigma_{3}\in\{0,1\}$ such that
$P(\sigma_{3}=0\mid\theta_{-3}=(\theta_{1},\theta_{2}))=P(\sigma_{3}=0\mid\theta_{-3}=(\theta_{1}^{\prime},\theta_{2}))=2/3$
and
$P(\sigma_{3}=1\mid\theta_{-3}=(\theta_{1},\theta_{2}^{\prime}))=P(\sigma_{3}=1\mid\theta_{-3}=(\theta_{1}^{\prime},\theta_{2}^{\prime}))=2/3,$
if accessed by agent 3, allows $F$ to satisfy Bayesian monotonicity.
Appendix B Appendix B: Omitted proofs
@@
Fix a agent $j.$ Since every agents’ beliefs and every signal has full support, the only possible signal that affects agent $j$’s expected payoff is $\sigma_{j}.$ Suppose $\sigma_{j}\in\{A_{j1},\dots,A_{jk}\}$ and let $q_{\ell}$ be the unconditional probability that $\sigma_{j}=A_{j\ell}$ (so $q_{\ell}$ is averaged over all possible states). If $F$ is incentive-compatible with respect to agent $j$ after agent $j$ receives this signal, then
@@
Rule 1: Suppose there is a SCR $F\in\widehat{F}$ and $i\in N$ such that every $m_{j}$ for $j\neq i$ is of the form (,,,)($\cdot$,$\varnothing$,$\cdot$,$\varnothing$)( ⋅ , ∅ , ⋅ , ∅ ). Additionally, $m_{i}$ is of the form (,,,)($\cdot$,$\varnothing$,$\cdot$,$\varnothing$)( ⋅ , ∅ , ⋅ , ∅ ), or (,,,).($\cdot$,$\cdot$,$\cdot$,$\varnothing$).( ⋅ , ⋅ , ⋅ , ∅ ) . Then $h(m)=F(\theta^{m}).$
Rule 2: Suppose there is a SCR $F\in\widehat{F}$ and $i\in N$ such that every $m_{j}$ for $j\neq i$ is of the form (,,,)($\cdot$,$\varnothing$,$\cdot$,$\varnothing$)( ⋅ , ∅ , ⋅ , ∅ ). Additionally, $m_{i}$ is of the form $(\cdot,\cdot,\cdot,y)$ for some $y\in X$. Let $\theta_{i}$ be the first entry of $m_{i},$ and define $y_{\theta_{i}}\in X$ by $y_{\theta_{i}}(\theta)=y(\theta_{-i},\theta_{i})$ for all $\theta\in\Theta.$ Then
@@
for $j=1,3$ and all $\theta_{i}^{\prime}\in\Theta_{i},$ and the inequalities are strict for type $\theta_{i}^{\prime}=\theta_{i}.$ Let $\sigma_{i}\in\{0,1\}$ be a binary signal such that if $\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4}\}\cup\alpha_{-i}^{-1}(\theta_{-i}^{3})$, then $P(\sigma_{i}=0\mid\theta_{-i})=\tau$ for some accuracy $10.5<\tau<$ 1. Additionally, if $\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4}\}\cup\alpha_{-i}^{-1}(\theta_{-i}^{3})$, then $P(\sigma_{i}=1\mid\theta_{-i})=\tau.$ With slight abuse of notation we write $\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}$ to mean $\{\theta_{-i}^{1},\theta_{-i}^{4}\}\cup\alpha_{-i}^{-1}(\theta_{-i}^{3}).$ Since agent $i$ can potentially undermine $\alpha$, we have from the definition that $\{\alpha_{-i}^{-1}(\theta_{-i}^{1}),\theta_{-i}^{2},\theta_{-i}^{3}\}\subset\Theta_{-i}\setminus\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}$. Thus, upon receiving $\sigma_{i}=0$, agent $i$ will place probability
@@
If the signal realization is $\sigma_{i}=1,$ one can similarly define $y:\Theta_{-i}\to A$ by $y(\theta_{-i}^{3})=F(\theta_{-i}^{4},\alpha_{i}(\theta_{i})),y(\theta_{-i}^{4})=F(\theta_{-i}^{3},\alpha_{i}(\theta_{i}))$, and $y(\theta_{-i})=F(\theta_{-i},\alpha_{i}(\theta_{i}))$ for all $\theta_{-i}\neq\theta_{-i}^{3},\theta_{-i}^{4}.$ The proof that $\alpha$ is undermined in this case is largely analogous. To show that
$\mathbb{E}(u_{i}(F(\theta_{-i},\theta_{i}^{\prime}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime},\sigma_{i}=1)\geq\mathbb{E}(u_{i}(y(\theta_{-i}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime},\sigma_{i}=1)$
for all $\theta_{i}^{\prime}\in\Theta_{i}$, we can use similar steps as above to show that $\tau$ needs to satisfy
$\tau\mu(\theta_{-i}^{3}\mid\theta_{i}^{\prime},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})\geq(1-\tau)\mu(\theta_{-i}^{4}\mid\theta_{i}^{\prime},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}),$
for all $\theta_{i}^{\prime}\in\Theta_{i}.$
Finally, one again needs to check that there is sufficiently large $\tau$ such that
$\mathbb{E}(u_{i}(F(\alpha(\theta_{-i},\theta_{i})),\theta_{i})\mid\theta_{i},\sigma_{i}=1)<\mathbb{E}(u_{i}(y(\alpha_{-i}(\theta_{-i})),\theta_{i})\mid\theta_{i},\sigma_{i}=1).$
Observe that
$\mathbb{E}(u_{i}$ $(F(\alpha(\theta_{-i},\theta_{i})),\theta_{i})\mid\theta_{i},\sigma_{i}=1)=(1-\tau)\mu(\theta_{-i}^{1}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{2},\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\mu(\theta_{-i}^{4}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\alpha_{-i}(\theta_{-i}^{4}),\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{3})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{3},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{4})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{4},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\begin{subarray}{c}\theta_{-i}^{\prime}\notin\alpha_{-i}^{-1}(\theta_{-i}^{4})\\ \theta_{-i}^{\prime}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}\end{subarray}}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\alpha_{-i}(\theta_{-i}^{\prime}),\alpha_{i}(\theta_{i})),\theta_{i}),$
while
$\mathbb{E}(u_{i}$ $(y(\alpha(\theta_{-i})),\theta_{i})\mid\theta_{i},\sigma_{i}=1)=(1-\tau)\mu(\theta_{-i}^{1}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{2},\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\mu(\theta_{-i}^{4}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\alpha_{-i}(\theta_{-i}^{4}),\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{3})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{4},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{4})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{3},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\begin{subarray}{c}\theta_{-i}^{\prime}\notin\alpha_{-i}^{-1}(\theta_{-i}^{4})\\ \theta_{-i}^{\prime}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}\end{subarray}}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\alpha_{-i}(\theta_{-i}^{\prime}),\alpha_{i}(\theta_{i})),\theta_{i}).$
It therefore suffices to pick $\tau$ sufficiently large so that
$(1-\tau)\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{3})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{3},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{4})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{4},\alpha_{i}(\theta_{i})),\theta_{i})$
$<(1-\tau)\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{3})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{4},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{4})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{3},\alpha_{i}(\theta_{i})),\theta_{i}).$
As before, this can be done with sufficiently large $\tau$. Thus, after agent $i$ receives the signal $\sigma_{i}$, she can undermine $\alpha$, and the proof is complete. ∎
@@
When $i\notin\{j_{1},\dots,j_{k}\}$ the signal $\sigma_{i}$ is uninformative so the inequalities don’t change between (v) and (vi). If $i\in\{j_{1},\dots,j_{k}\}$, then truth-telling is a dominant strategy for agent $i$, so it must be that
@@
for all $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$, since $\mu$ and $\sigma_{i}$ have full support. Thus both (v) and (vi) are imply

### v4

@@ 6 Conclusion
The main finding of the paper is that increasing the information agents have about the types of other agents can serve to eliminate undesired equilibria. We provide a sufficient condition under which a planner interested in full implementation will prefer agents have information about other agents beyond the common prior. The condition is straightforward to verify, as it requires only some strict preferences from some agents over type profiles of other agents and does not rely on any specific payoff structures. A future direction can be to relax the full-support signal assumption. Doing so would necessitate careful modeling of type spaces but could prove even more useful for the planner: Such signals can immediately enlarge the strategy spaces for agents in the direct mechanism, which can not only eliminate undesired equilibria but also cure violations of incentive compatibility.
Appendix A Appendix: Omitted proofs
@@
Fix an agent $j.$ Since every agents’ beliefs and every signal has full support, the only possible signal that affects agent $j$’s expected payoff is $\sigma_{j}.$ Suppose $\sigma_{j}\in\{A_{j1},\dots,A_{jk}\}$ and let $q_{\ell}$ be the unconditional probability that $\sigma_{j}=A_{j\ell}$ (so $q_{\ell}$ is averaged over all possible states). If $F$ is incentive-compatible with respect to agent $j$ after agent $j$ receives this signal, then
@@
Rule 1: Suppose there is an $i\in I$ such that every $m_{j}$ for $j\neq i$ is of the form (,,,)($\cdot$,$\varnothing$,$\cdot$,$\varnothing$)( ⋅ , ∅ , ⋅ , ∅ ). Additionally, $m_{i}$ is of the form (,,,)($\cdot$,$\varnothing$,$\cdot$,$\varnothing$)( ⋅ , ∅ , ⋅ , ∅ ), or (,,,).($\cdot$,$\cdot$,$\cdot$,$\varnothing$).( ⋅ , ⋅ , ⋅ , ∅ ) . Then $h(m)=F(\theta^{m}).$
Rule 2: Suppose there is an $i\in I$ such that every $m_{j}$ for $j\neq i$ is of the form (,,,)($\cdot$,$\varnothing$,$\cdot$,$\varnothing$)( ⋅ , ∅ , ⋅ , ∅ ). Additionally, $m_{i}$ is of the form $(\cdot,\cdot,\cdot,y)$ for some $y\in X$. Let $\theta_{i}$ be the first entry of $m_{i},$ and define $y_{\theta_{i}}\in X$ by $y_{\theta_{i}}(\theta)=y(\theta_{-i},\theta_{i})$ for all $\theta\in\Theta.$ Then
@@
for $j=1,3$ and all $\theta_{i}^{\prime}\in\Theta_{i},$ and each inequality is strict for some type $\theta_{i}.$ Let $\sigma_{i}\in\{0,1\}$ be a binary signal such that if $\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4}\}\cup\alpha_{-i}^{-1}(\theta_{-i}^{3})$, then $P(\sigma_{i}=0\mid\theta_{-i})=\tau$ for some accuracy $10.5<\tau<$ 1. Additionally, if $\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4}\}\cup\alpha_{-i}^{-1}(\theta_{-i}^{3})$, then $P(\sigma_{i}=1\mid\theta_{-i})=\tau.$ With slight abuse of notation we write $\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}$ to mean $\{\theta_{-i}^{1},\theta_{-i}^{4}\}\cup\alpha_{-i}^{-1}(\theta_{-i}^{3}).$ Since agent $i$ can potentially undermine $\alpha$, we have from the definition that $\{\alpha_{-i}^{-1}(\theta_{-i}^{1}),\theta_{-i}^{2},\theta_{-i}^{3}\}\subset\Theta_{-i}\setminus\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}$. Thus, upon receiving $\sigma_{i}=0$, agent $i$ will place probability
@@
The proof is analogous for the realization $\sigma_{i}=1.$
@@
When $i\notin\{j_{1},\dots,j_{k}\}$ the signal $\sigma_{i}$ is uninformative so the inequalities don’t change between (1) and (2). If $i\in\{j_{1},\dots,j_{k}\}$, then truth-telling is a dominant strategy for agent $i$, so it must be that
@@
for all $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$, since $\mu$ and $\sigma_{i}$ have full support. Thus both (1) and (2) imply